OpenStudy (anonymous):

Determine all values of c so that the lines cy+5x=7 and cy-3x=6 are perpendicular.

OpenStudy (anonymous):

okay, let's switch these both into y=mx+b form, because its a little easier to visualize slope that way. we get the first line: y=-(5/c)x+7/c, and the 2nd line we get y=(3/c)x+6/c. the only qualification for two lines to be perpendicular is that the slope of one is equal to the negative reciprocal of the other (the reciprocal of n is (-1/n) for example) so, (-5/c) must be equal to -1/(3/c). we simplify both and we get -5/c = -c/3. we multiply both sides by c and get -5=-(c^2)/3 multiply both sides by -3 and we get that 15=c^2 so c= +/- \[\sqrt{15}\]

OpenStudy (anonymous):

@JHall96 Thank you so much!